OpenStudy (anonymous):

I want to determine the first derivatives. y=(x+2)^2(2x-1)^3

OpenStudy (anonymous):

the solution step by step

OpenStudy (australopithecus):

YOu dont need to use the definition of a derivative do you?

OpenStudy (anonymous):

I cannot understand all of above at all. Do you have any easier method than this.

OpenStudy (australopithecus):

this is how I do derivatives, y=(x+2)^2(2x-1)^3 Split it into smaller functions, g(x) = (x+2)^2 m(x) =(2x-1)^3 We can split these new functions into even simpler functions so lets do that Starting with g(x) d(x) = x^2 s(x) = x+2 now take the derivative of these two smaller functions, d(x) = x^2 d'(x) = 2x s(x) = x + 2 s'(x) = 1 Now we notice that this is one function in another function so we can just apply chain rule which is, g'(x) = d'(s(x))*s'(x) sub in, g(x) = (x+2)^2 g'(x) = 2(x+2)*1 Now do the same with m(x) you should notice right away that you have to use product rule with g(x) and m(x) so just use the product rule formula which is, f'(x) = g'(x)m(x) + g(x)m'(x)

OpenStudy (australopithecus):

Follow my method

OpenStudy (australopithecus):

it is the best way to take a derivative you will never make a mistake and you will train your brain to be able to solve them in your head

OpenStudy (australopithecus):

if you have any questions please ask, I can complete the problem if you are confused

OpenStudy (australopithecus):

I can walk you through m(x) and the final derivative if you like but I would prefer that you tried my method first

OpenStudy (australopithecus):

try deriving m(x) then the final derivative, I have no idea why this isn't taught in most schools

OpenStudy (anonymous):

wahh, thanks a lot Australopithecus but... I have a problem with complete this using f'(x) = g'(x)m(x) + g(x)m'(x) Nt very expert. I find it little hard

OpenStudy (australopithecus):

you understand function notation right? that g(x) is just another way of writing (x+2)^2 it is just faster

OpenStudy (australopithecus):

Can you show me what you got for the derivative of m(x)? Just do it exactly as how I did the derivative of g(x) If you dont have the basic rules of derivatives down then I recommend having them in front of you

OpenStudy (australopithecus):

Once you get this method down you only really need to memorize chain rule, product rule and maybe quaint rule depending on if your prof cares

OpenStudy (anonymous):

for m(x) = (2x-1)^3 m'(x) = 6(2x-1)^2 correct?

OpenStudy (australopithecus):

looks good to me

OpenStudy (australopithecus):

now just sub into the product rule equation

OpenStudy (australopithecus):

g(x) and m(x) its that easy

OpenStudy (australopithecus):

BTW since I showed you chain rule and product rule. I might as well show you quotient rule which is: f(x) = g(x)/s(x) f'(x) = [g'(x)s(x) - s'(x)g(x)]/(s(x))^2

OpenStudy (australopithecus):

those are the three formulas you have to memorize

OpenStudy (australopithecus):

practice using my method and this stuff will be a joke to you

OpenStudy (anonymous):

ok, i will . Here what I got this after subs it, 2(x+2)(2x-1)^3 + (x+2)^2 (6(2x-1))^2 But i dont know how to complete this. can you show please...

OpenStudy (australopithecus):

tbh though quotient rule is useless it is just derived from product rule, simply use the rule: \[x^{-1} = \frac{1}{x}\] and you can use product rule always, for example f(x) = s(x)/d(x) f(x) = s(x)d(x)^-1 f'(x) = s'(x)d(x) + s(x)d'(x)^-1

OpenStudy (australopithecus):

2(x+2)(2x-1)^3 + (x+2)^2 (6(2x-1))^2 This is wrong, look at it and you will see why right away

OpenStudy (australopithecus):

its not that you didnt apply the method correctly

OpenStudy (australopithecus):

you made an algebraic mistake

OpenStudy (anonymous):

which one, i cant see it

OpenStudy (australopithecus):

2(x+2)(2x-1)^3 + (x+2)^2 (6(2x-1))^2 /\

OpenStudy (australopithecus):

see it now?

OpenStudy (australopithecus):

you squared the 6

OpenStudy (australopithecus):

(6(2x-1))^2 = 6^2(2x-1))^2 =/= 6(2x-1)^2

OpenStudy (australopithecus):

(6(2x-1))^2 = 6^2(2x-1))^2 =/= 6(2x-1)^2 = m'(x)

OpenStudy (australopithecus):

show me the answer :D

OpenStudy (australopithecus):

careful when subbing into the formulas

OpenStudy (anonymous):

2(x+2)(2x-1)^3 + (x+2)^2 (6(2x-1)^2) like this?

OpenStudy (anonymous):

then help me how to complete this

OpenStudy (anonymous):

I would just expand out the polynomial and then take the derivative.

OpenStudy (australopithecus):

you cant avoid chain rule

OpenStudy (australopithecus):

that is the answer

OpenStudy (anonymous):

but not the final answer... how must I do it?

OpenStudy (australopithecus):

f'(x) = g'(x)m(x) + g(x)m'(x) m(x) = (2x-1)^3 m'(x) = 6(2x-1)^2 g(x) = (x+2)^2 g'(x) = 2(x+2)*1 f'(x) = 2(x+2)(2x-1)^3 + 6(2x-1)^2((x+2)^2)

OpenStudy (australopithecus):

that is the final answer

OpenStudy (australopithecus):

they might want you to simplify it

OpenStudy (australopithecus):

but that is the derivative

OpenStudy (anonymous):

yup, tht right

OpenStudy (australopithecus):

be very very careful how you set up your brakets

OpenStudy (australopithecus):

you would have been right but you placed your brackets incorrectly and squared the 6 in m'(x)

OpenStudy (anonymous):

the final answr suppose to be: 10(2x-1)^2(x+2)(x+1) how is it done? show me

OpenStudy (australopithecus):

@wio that is awful advice what if she was asked to take the derivative of f(x) = (3 + x)^99 You can't avoid chain rule in calculus

OpenStudy (australopithecus):

just simplify

zepdrix (zepdrix):

Your first question?\[\Large \color{royalblue}{\text{Welcome to OpenStudy Sytee! :)}}\]

OpenStudy (anonymous):

@zepdrix yup, thanks :')

OpenStudy (anonymous):

don know how to simplify it

OpenStudy (usukidoll):

what's up @zepdrix

OpenStudy (australopithecus):

f'(x) = 2(x+2)(2x-1)^3 + 6(2x-1)^2((x+2)^2) collect like terms first off f'(x) = (x+2)(2x-1)^2[2(2x-1) + 6(x+2)]

OpenStudy (australopithecus):

can you do the next step? I'm simply using the rule a(x + b) = ax + ab

OpenStudy (australopithecus):

in the last step

OpenStudy (australopithecus):

you are almost to where you want to be with, f'(x) = (x+2)(2x-1)^2[2(2x-1) + 6(x+2)] Just fool around with, 2(2x-1) + 6(x+2)

OpenStudy (australopithecus):

You seriously need to have algebra mastered if you are going to be successful in calculus. I recommend looking up math rules and committing them to memory if you have any doubts about what you can and cannot do mess around with, www.wolframalpha.com to convince yourself of the basic math rules

OpenStudy (anonymous):

ok, i get it now. thankyou

OpenStudy (australopithecus):

ok awesome :D if you have any other questions feel free to message me take care and good luck

OpenStudy (anonymous):

i have a lot of question to ask you actually

OpenStudy (australopithecus):

oh really

OpenStudy (australopithecus):

well Im going to sleep soon so if you have them ask them

OpenStudy (anonymous):

um, if you don mind, show me the next step plz

OpenStudy (australopithecus):

can you at least expand them 2(2x-1) + 6(x+2)

OpenStudy (anonymous):

(4x-2)+(6x+12)

OpenStudy (anonymous):

if you give me all the solution step, maybe I can revise it and look bck what I really didnt understand it.

OpenStudy (australopithecus):

you can remove the brakets

OpenStudy (australopithecus):

I shouldnt have to its pretty much solved

OpenStudy (australopithecus):

you can do this algebra I believe in you!

OpenStudy (australopithecus):

Add/subtract like terms and use a(x+b) = ax + ab and you are done

OpenStudy (anonymous):

ok, i will try now..

OpenStudy (australopithecus):

ooops

OpenStudy (australopithecus):

sorry it should be 4x-2 + 6x+12 so you have, (x+2)(2x-1)^2[4x-2 + 6x+12] and you want 10(2x-1)^2(x+2)(x+1)

OpenStudy (australopithecus):

do you have it solved?

OpenStudy (anonymous):

not yet, still confusing me. this one, (x+2)(2x-1)^2 must I multiple each of them

OpenStudy (australopithecus):

no they are in your final answer leave them alone

OpenStudy (australopithecus):

you only care about [4x-2 + 6x+12]

OpenStudy (australopithecus):

or you should only care about simplifying that

OpenStudy (anonymous):

omg, i solve it!

OpenStudy (australopithecus):

good job

OpenStudy (anonymous):

thanks a lot!!!

OpenStudy (australopithecus):

my pleasure

OpenStudy (anonymous):

you may go to ur beauty sleep now. :)

OpenStudy (australopithecus):

Ok good to hear you are done with me, good luck with calc